# Interchangeable Token Embeddings for Extendable Vocabulary and Alpha-Equivalence

### v2

@@ Abstract
We propose a novel approach for learning interchangeable tokens in language models to obtain an extendable vocabulary that can generalize to new tokens. Our method addresses alpha-equivalence, the principle that renaming bound variables preserves semantics. This property arises in many formal languages such as temporal logics, where all proposition symbols represent the same concept but remain distinct. To handle such tokens, we develop a dual-part embedding approach. The first part is shared across all interchangeable tokens, enforcing that they represent the same core concept. The second part is randomly generated for each token, enabling distinguishability. As a baseline, we consider a simpler approach that uses alpha-renaming for data augmentation. We also present alpha-covariance, a metric for measuring robustness against alpha-conversions. When evaluated in a Transformer encoder-decoder model for solving linear temporal logic formulae and copying with extendable vocabulary, our method demonstrates promising generalization capabilities as well as a favorable inductive bias for alpha-equivalence.
@@ 1 Introduction
Following the deep learning revolution that affected numerous application areas (Dargan et al., 2020), recent literature shows that deep learning based approaches also perform well in neurosymbolic reasoning tasks, such as theorem proving (Han et al., 2021) and mathematical reasoning (Rabe et al., 2020). The formal reasoning capabilities of these models were once doubted, but Liu et al. (2023) demonstrated the ability of Transformer models (Vaswani et al., 2017) to learn shortcuts to automata. Of particular interest is the generalization ability of such models to unseen, out-of-distribution data (Sanh et al., 2021), enhancing their appeal for logical reasoning (Abbe et al., 2023).
Another application area is linear-time temporal logic (LTL), which is heavily utilized by the formal verification community (Clarke et al., 2018; Baier and Katoen, 2008) for reasoning about how logical propositions change over time (Pnueli, 1977). Through the use of temporal operators, LTL formulae can specify, for example, that a proposition $p$ must hold at all time steps ($\mathbf{G}p$), or at least one time step ($\mathbf{F}p$). LTL formulae operate on traces, which describe how the propositions change over time.
Solving a given LTL formula involves finding a satisfying trace, and it proved essential for generating examples for system specifications in the literature. This field was dominated by the methods that use classical algorithms, such as spot (Duret-Lutz et al., 2022) and aalta (Li et al., 2014). However, following the success of Transformer models on end-to-end symbolic integration (Lample and Charton, 2019), Hahn et al. (2021) attacked the LTL solving problem using the same approach. Their capability to generalize to longer formulae is especially noteworthy, and it was made possible thanks to tree-positional encoding (Shiv and Quirk, 2019).
@@
In this paper, we propose a novel approach for representing interchangeable tokens in neural network models. To summarize, our method constructs some part of the token embeddings on-the-fly instead of learning all of them during training. The token embeddings for interchangeable tokens consist of two parts: a learnable part and a randomized part. The learnable part is shared across all interchangeable tokens, and the model must depend on the randomized part to differentiate these tokens. We use the weight tying technique (Press and Wolf, 2016) to share the same token embeddings with the final projection matrix, which calculates the logits (i.e., next-token probabilities before softmax).
We use our embedding method in a Transformer encoder-decoder model and evaluate it on two tasks: copying with an extendable vocabulary and solving LTL formulae. For the second task, we use datasets generated using spot (Duret-Lutz et al., 2022), a library for LTL manipulation and model checking using conventional algorithms. As a baseline, we consider a simpler approach that uses alpha-renaming for data augmentation during training to expose the model to a larger vocabulary, which is also new in the literature to the best of our knowledge. Overall, our method demonstrates generalization capabilities to larger vocabulary sizes, and also combines well with positional encodings that exhibit length generalization. We also experiment with dataset perturbation to show that our method introduces a helpful inductive bias for alpha-equivalence. Finally, we present alpha-covariance, a metric for measuring robustness against alpha-conversions.
@@ 2 Related Work
Language modeling and formal reasoning
The transformer architecture (Vaswani et al., 2017), now ubiquitous in modern deep learning, was initially proposed as a generative model to translate between natural languages autoregressively. This led to many successful attempts to frame formal reasoning tasks as language modeling problems, such as symbolic integration (Lample and Charton, 2019), symbolic regression (Kamienny et al., 2022; Vastl et al., 2022), LTL solving (Hahn et al., 2021), and many more. Further developments shifted the field towards large language models (LLMs), e.g., by prompting a model pre-trained on a gigantic scale (Frieder et al., 2023), by enhancing the prompt with retrieved references for proof generation (Welleck et al., 2022; Yang et al., 2023), by training an LLM on a specialized dataset for mathematics (Azerbayev et al., 2023). However, the reasoning abilities of LLMs were questioned by (Tang et al., 2023), who showed LLMs struggle with symbolic reasoning when semantics are decoupled, and by others (Wu et al., 2023).
Extensible vocabulary
Efforts to create an extensible vocabulary for neural networks are scarce in the broader machine learning community, let alone the formal reasoning literature. Morazzoni et al. (2023) exploited dictionary definitions to create extensible word embeddings. Wei et al. (2016) proposed a vocabulary-extensible sign language recognition framework by using a component based approach, where each sign gesture is recognized based on common components such as hand shape, orientation, axis, rotation, and trajectory. These studies depend on either external information (dictionary definitions) or properties specific to an application area (components of hand gesture); they do not attempt to design an extensible vocabulary for interchangeable tokens, which has been neglected by the literature alongside the concept of alpha-equivalence.
@@ 3 Problem Definition
In a language modeling problem, the goal is to predict the next token in the output sequence given the input and the past output. (See Appendix A.1 for more details on language models.) Let ${\mathbb{V}}$ denote the set of all unique tokens, i.e., the vocabulary of a language modeling problem. We assume that ${\mathbb{V}}_{i}$ is the set of interchangeable tokens and ${\mathbb{V}}_{n}={\mathbb{V}}\backslash{\mathbb{V}}_{i}$ is the set of non-interchangeable tokens. The core idea behind alpha-equivalence is that renaming interchangeable tokens between each other in both input and output preserves meaning. Let $f\colon{\mathbb{V}}\to{\mathbb{V}}$ be a bijection such that $f(x)=x$ for all $x\in{\mathbb{V}}_{n}$, i.e., $f$ arbitrarily renames the interchangeable tokens between each other in one-to-one correspondence and preserves the rest of the tokens. We apply $f$ to each token in a given pair of input ${\bm{a}}\in{\mathbb{V}}^{*}$ and output ${\bm{b}}\in{\mathbb{V}}^{*}$ strings, obtaining ${\bm{a}}^{\prime}=(f({a}_{1}),f({a}_{2}),\ldots)$ and ${\bm{b}}^{\prime}=(f({b}_{1}),f({b}_{2}),\ldots)$. We call this operation alpha-conversion or alpha-renaming. The set of interchangeable tokens ${\mathbb{V}}_{i}$ must be defined such that ${\bm{a}}^{\prime}$ and ${\bm{b}}^{\prime}$ form a valid input-output pair semantically equivalent to $({\bm{a}},{\bm{b}})$ for all possible $f$.
@@
Example
In the LTL solving problem (Appendix A.2), the set of non-interchangeable tokens ${\mathbb{V}}_{n}$ includes the operators, constants, delimiter tokens (“;”, “{”, “}”), and any special tokens such as the end token. The set of interchangeable tokens equals to the set of atomic propositions (APs): ${\mathbb{V}}_{i}=P$. Assuming $P=\{\texttt{a},\texttt{b}\}$, the formula-trace pair (“&aXb”, “a;b;{1}”) is alpha-equivalent to (“&bXa”, “b;a;{1}”). Further, assume that the augmented set of interchangeable tokens is ${\mathbb{V}}_{i}^{\prime}=P^{\prime}=\{\texttt{a},\texttt{b},\texttt{c},\texttt{d}\}$. Now, the aforementioned pair can also be equivalently represented as (“&cXd”, “c;d;{1}”). The augmented vocabulary allows the expression of formula-trace pairs that feature up to 4 APs instead of 2. For example, (“&&abX&cd”, “&ab;&cd;{1}”) cannot be expressed using $P=\{\texttt{a},\texttt{b}\}$. Our goal is to create a model that can handle such inputs despite being trained on the limited vocabulary ${\mathbb{V}}={\mathbb{V}}_{n}\cup P$.
@@ 4.1 Embedding matrix
Construction of the embedding matrix
For a vocabulary with $n$ non-interchangeable tokens and $m$ interchangeable tokens, ${\bm{L}}\in{\mathbb{R}}^{n\times d_{\alpha}}$ represents the matrix of learnable embeddings for non-interchangeable tokens, ${\bm{\alpha}}\in{\mathbb{R}}^{1\times d_{\alpha}}$ the shared learnable embedding for interchangeable tokens, and ${\bm{\beta}}_{i}\in{\mathbb{R}}^{1\times d_{\beta}}$ the randomly-generated embedding for the $i$th interchangeable token where $1\leq i\leq m$. Note that ${\bm{\alpha}}$ and ${\bm{\beta}}_{i}$ are row vectors. A zero matrix of size $i\times j$ is represented by $\bm{0}^{i,j}$. In addition, we define two row-based L2 normalization functions $f_{bn}({\bm{X}})$ and $f_{fn}({\bm{X}})$ that divide each row ${\bm{X}}_{i,:}$ by its L2 norm $||{\bm{X}}_{i,:}||$. These two functions are identical but can be disabled independently from each other, hence the separation. Finally, the overall structure of the embedding matrix ${\bm{U}}$ is shown in Equation 1. In this construction, the interchangeable tokens are assumed to come after the non-interchangeable tokens. Note that it’s also possible to implement multiple sets of different interchangeable tokens via a trivial extension.
@@
Normalization
There are several concerns that warrant the heavy use of normalization while constructing ${\bm{U}}$, as seen in Equation 1. Firstly, $d_{\alpha}$ dimensions and $d_{\beta}$ dimensions should not overwhelm each other in terms of magnitude. Normalizing ${\bm{\alpha}}$ and ${\bm{\beta}}_{i}$ separately addresses this issue. The magnitude of the concatenated embedding is another concern, which is handled by the final normalization. The normalization of ${\bm{L}}$ is redundant (since the final normalization does the same operation after the concatenation with zeros) but kept in Equation 1 for readability.
@@ 4.2 Random embedding generation
Table 1: Comparison of random vector generation methods.
Method Normal Distribution Neighboring Points Hypercube Vertices
Formula ${\textnormal{a}}_{i}\sim\mathcal{N}(0,1)$ ${\textnormal{a}}_{i}\in\{-1,0,1\}$ ${\textnormal{a}}_{i}\in\{-1,1\}$
$|{\mathbf{a}}_{i}|\neq 0$
Size for $n$-dims Continuous $13^{n}-$ 1 $2^{n}$
Sample Visualization [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
In the normal distribution method, we don’t have any additional constraints to ensure distinguishability between vectors. However, in other two methods, we need make sure that each interchangeable token gets assigned to a unique vector since the sampling set is finite. To achieve this quickly and space-efficiently, we define a mapping from integers to possible vectors. The unique vectors are generated by sampling $m$ unique random integers (which can be calculated efficiently using the reservoir sampling technique), and then using the defined mapping to convert these integers to the vectors. This strategy avoids materializing the whole set of possible vectors. In the hypercube vertices method, we map the binary digits of an integer in $[0,2^{d_{\beta}})$ to $\{-1,1\}$. Although “Neighboring Points" is simply the ternary version of the same idea, avoiding the zero vector requires special care. The zero vector maps to the integer $i_{z}=(3^{d_{\beta}}-1)/2$. Therefore, we define our domain as the integers in $[0,3^{d_{\beta}}-1)$ and add $1$ to the integer $i$ before converting it if $i\geq i_{z}$.
Integer mapping approach for generating unique vectors works well for up to 32 dimensions, after which the limits of integer representation become an issue for reservoir sampling. Therefore, in such cases, we simply disable the uniqueness check because the size of the sampling set grows exponentially, rendering the probability of drawing the same sample negligible.
@@ 4.3 Projection
In a traditional language modeling setting, since both the embedding and projection matrices are entirely composed of learnable parameters, it’s not necessary to share them, even though there are many advantages of weight tying (Press and Wolf, 2016). However, we construct the embedding matrix manually in our method, which makes weight tying a requirement. Furthermore, since we perform our experiments on an encoder-decoder architecture in this paper, we utilize a three-way weight tying approach, whereby the embedding matrices of encoder and decoder are tied in addition to the final projection matrix. Three-way weight tying is particularly appropriate for our problem domain since many tokens are shared between the LTL formulae and traces.
Feature normalization
@@
Given the output of the last layer before the final projection ${\bm{v}}$ (henceforth called feature vector), instead of directly applying the final projection as in ${\bm{U}}{\bm{v}}$, we apply L2 normalization to the feature vector ${\bm{v}}$ before passing it through the final projection: ${\bm{U}}f_{fn}({\bm{v}})$. This matrix multiplication constitutes taking a dot product with each row. Since ${\bm{a}}\cdot{\bm{b}}=|{\bm{a}}||{\bm{b}}|cos(\theta)$ where $\theta$ is the angle between ${\bm{a}}$ and ${\bm{b}}$, normalizing both the embeddings and the feature vector leaves only the cosine term to determine the logits. This forces the model to distinguish between tokens based solely on the directions, which may improve the gradient flow.
Cosine loss
@@
If we normalize both the embeddings and the feature vector, the only thing that determines each logit is the cosine of the angle between the feature vector and the embedding. Applying the softmax loss to such logits is known as cosine loss in the literature. Although cosine-based loss functions were successful in face recognition (Ranjan et al., 2017; Wang et al., 2017), it proved sensitive to hyperparameter settings in these losses. To avoid this problem, we use AdaCos loss function (Zhang et al., 2019) that scales the logits adaptively throughout training.
@@
To adapt the AdaCos loss function to our use case, we make the following modifications: Since the language modeling problem involves a sequence length dimension in addition to the batch dimension, we combine these two dimensions while ignoring the padding tokens, effectively treating both dimensions as batch dimensions. However, since this change greatly increases the number of batch dimensions, it can lead to numerical issues, even with the log-sum-exp trick. Therefore, we clip the scale value calculated by AdaCos to a maximum of 100 to avoid numerical issues.

### v3

@@ Abstract
Language models lack the notion of interchangeable tokens: symbols that are semantically equivalent yet distinct, such as bound variables in formal logic. This limitation prevents generalization to larger vocabularies and hinders the model’s ability to recognize alpha-equivalence, where renaming bound variables preserves meaning. We formalize this machine learning problem and introduce alpha-covariance, a metric for evaluating robustness to such transformations. To tackle this task, we propose a dual-part token embedding strategy: a shared component ensures semantic consistency, while a randomized component maintains token distinguishability. Compared to a baseline that relies on alpha-renaming for data augmentation, our approach demonstrates improved generalization to unseen tokens in linear temporal logic solving, propositional logic assignment prediction, and copying with an extendable vocabulary, while introducing a favorable inductive bias for alpha-equivalence. Our findings establish a foundation for designing language models that can learn interchangeable token representations, a crucial step toward more flexible and systematic reasoning in formal domains. Our code and project page are available at necrashter.github.io/interchangeable-token-embeddings
Machine Learning, ICML
@@ 1 Introduction
Following the deep learning revolution that affected numerous application areas (Dargan et al., 2020), recent literature shows that deep learning based approaches also perform well in neurosymbolic reasoning tasks, such as theorem proving (Han et al., 2021) and mathematical reasoning (Rabe et al., 2020). The formal reasoning capabilities of these models were once doubted, but Liu et al. (2023) demonstrated the ability of Transformer models (Vaswani et al., 2017) to learn shortcuts to automata. Of particular interest is the generalization ability of such models to unseen, out-of-distribution data (Sanh et al., 2022), enhancing their appeal for logical reasoning (Abbe et al., 2023).
Another application area is linear-time temporal logic (LTL), which is heavily utilized by the formal verification community (Clarke et al., 2018; Baier & Katoen, 2008) for reasoning about how logical propositions change over time (Pnueli, 1977). Through the use of temporal operators, LTL formulae can specify, for example, that a proposition $p$ must hold at all time steps ($\mathbf{G}p$), or at least one time step ($\mathbf{F}p$). LTL formulae operate on traces, which describe how the propositions change over time.
Solving a given LTL formula involves finding a satisfying trace, and it proved essential for generating examples for system specifications in the literature. This field was dominated by the methods that use classical algorithms, such as spot (Duret-Lutz et al., 2022) and aalta (Li et al., 2014). However, following the success of Transformer models on end-to-end symbolic integration (Lample & Charton, 2019), Hahn et al. (2021) attacked the LTL solving problem using the same approach. Their capability to generalize to longer formulae is especially noteworthy, and it was made possible thanks to tree-positional encoding (Shiv & Quirk, 2019).
@@
In this paper, we propose a novel approach for representing interchangeable tokens in neural network models. To summarize, our method constructs some part of the token embeddings on-the-fly instead of learning all of them during training. The token embeddings for interchangeable tokens consist of two parts: a learnable part and a randomized part. The learnable part is shared across all interchangeable tokens, and the model must depend on the randomized part to differentiate these tokens. Thanks to the randomized component, our method can generate embeddings for arbitrarily many interchangeable tokens as needed during both training and inference, with the only practical limitation being the exponentially growing sampling set size for discrete random generation methods. We use the weight tying technique (Press & Wolf, 2016) to share the same token embeddings with the final projection matrix, which calculates the logits (i.e., next-token probabilities before softmax).
We use our embedding method in a Transformer encoder-decoder model and evaluate it on three tasks: copying with an extendable vocabulary, solving LTL formulae, and predicting assignments for propositional logic. As a baseline, we consider a simpler approach that uses alpha-renaming for data augmentation during training to expose the model to a larger vocabulary, which is also new in the literature to the best of our knowledge. Overall, our method demonstrates generalization capabilities to larger vocabulary sizes, and also combines well with positional encodings that exhibit length generalization. We also experiment with dataset perturbation to show that our method introduces a helpful inductive bias for alpha-equivalence. Finally, we present alpha-covariance, a metric for measuring robustness against alpha-conversions that is applicable to any domain where alpha-equivalence is relevant.
Overall, our contributions can be summarized as follows.
  1. 1.
    Identify the problem of generalizing to larger vocabularies in (formal) language modeling tasks, and define an experimental protocol to study this problem.
  2. 2.
    Propose alpha-covariance, a novel metric for measuring robustness against alpha-conversions, applicable to any domain with interchangeable tokens.
  3. 3.
    Introduce a dual-part embedding method for vocabulary generalization and improved alpha-covariance, with negligible computational overhead.
  4. 4.
    Verify the proposed method thoroughly on three tasks: copying with extendable vocabulary, solving LTL formulae, and predicting assignments for propositional logic.
@@ 2 Related Work
Language modeling and formal reasoning. The transformer architecture (Vaswani et al., 2017), now ubiquitous in modern deep learning, was initially proposed as a generative model to translate between natural languages autoregressively. This led to many successful attempts to frame formal reasoning tasks as language modeling problems, such as symbolic integration (Lample & Charton, 2019), symbolic regression (Kamienny et al., 2022; Vastl et al., 2022), LTL solving (Hahn et al., 2021), and many more. Further developments shifted the field towards large language models (LLMs), e.g., by prompting a model pre-trained on a gigantic scale (Frieder et al., 2023), by enhancing the prompt with retrieved references for proof generation (Welleck et al., 2022; Yang et al., 2023), by training an LLM on a specialized dataset for mathematics (Azerbayev et al., 2023). However, the reasoning abilities of LLMs were questioned by (Tang et al., 2023), who showed LLMs struggle with symbolic reasoning when semantics are decoupled, and by others (Wu et al., 2023).
Extensible vocabulary. Efforts to create an extensible vocabulary for neural networks are scarce in the broader machine learning community, let alone the formal reasoning literature. Morazzoni et al. (2023) exploited dictionary definitions to create extensible word embeddings. Wei et al. (2016) proposed a vocabulary-extensible sign language recognition framework by using a component based approach, where each sign gesture is recognized based on common components such as hand shape, orientation, axis, rotation, and trajectory. These studies depend on either external information (dictionary definitions) or properties specific to an application area (components of hand gesture); they do not attempt to design an extensible vocabulary for interchangeable tokens, which has been neglected by the literature alongside the concept of alpha-equivalence.
@@ 3 Problem Definition
In language modeling, the goal is to predict the next token in the output sequence given the input and the past output. (See Appendix A for more background.) Let ${\mathbb{V}}$ denote the set of all unique tokens, i.e., the vocabulary of a language modeling problem. We use ${\mathbb{V}}^{*}$ to denote the set of all finite sequences of tokens (strings) from ${\mathbb{V}}$. We assume that ${\mathbb{V}}_{i}$ is the set of interchangeable tokens and ${\mathbb{V}}_{n}={\mathbb{V}}\backslash{\mathbb{V}}_{i}$ is the set of non-interchangeable tokens. The core idea behind alpha-equivalence is that renaming interchangeable tokens between each other in both input and output preserves meaning. Let $f\colon{\mathbb{V}}\to{\mathbb{V}}$ be a bijection such that $f(x)=x$ for all $x\in{\mathbb{V}}_{n}$, i.e., $f$ arbitrarily renames the interchangeable tokens between each other in one-to-one correspondence and preserves the rest of the tokens. We apply $f$ to each token in a given pair of input ${\bm{a}}\in{\mathbb{V}}^{*}$ and output ${\bm{b}}\in{\mathbb{V}}^{*}$ strings, obtaining ${\bm{a}}^{\prime}=(f({a}_{1}),f({a}_{2}),\ldots)$ and ${\bm{b}}^{\prime}=(f({b}_{1}),f({b}_{2}),\ldots)$. We call this operation alpha-conversion or alpha-renaming. The set of interchangeable tokens ${\mathbb{V}}_{i}$ must be defined such that ${\bm{a}}^{\prime}$ and ${\bm{b}}^{\prime}$ form a valid input-output pair semantically equivalent to $({\bm{a}},{\bm{b}})$ for all possible $f$.
@@
Example. In the LTL solving problem (Appendix B), the set of non-interchangeable tokens ${\mathbb{V}}_{n}$ includes the operators, constants, delimiter tokens (“;”, “{”, “}”), and any special tokens such as the end token. The set of interchangeable tokens equals to the set of atomic propositions (APs): ${\mathbb{V}}_{i}=P$. Assuming $P=\{\texttt{a},\texttt{b}\}$, the formula-trace pair (“&aXb”, “a;b;{1}”) is alpha-equivalent to (“&bXa”, “b;a;{1}”). Further, assume that the augmented set of interchangeable tokens is ${\mathbb{V}}_{i}^{\prime}=P^{\prime}=\{\texttt{a},\texttt{b},\texttt{c},\texttt{d}\}$. Now, the aforementioned pair can also be equivalently represented as (“&cXd”, “c;d;{1}”). The augmented vocabulary allows the expression of formula-trace pairs that feature up to 4 APs instead of 2. For example, (“&&abX&cd”, “&ab;&cd;{1}”) cannot be expressed using $P=\{\texttt{a},\texttt{b}\}$. Our goal is to create a model that can handle such inputs despite being trained on the limited vocabulary ${\mathbb{V}}={\mathbb{V}}_{n}\cup P$.
@@ 4 Proposed Method
[…]
tableComparison of random vector generation methods. Method Normal Distribution Neighboring Points Hypercube Vertices Formula ${\textnormal{a}}_{i}\sim\mathcal{N}(0,1)$ ${\textnormal{a}}_{i}\in\{-1,0,1\}$ ${\textnormal{a}}_{i}\in\{-1,1\}$ $\|{\mathbf{a}}\|\neq 0$ Size for $n$-dims Continuous $13^{n}-$ 1 $2^{n}$ Sample Visualization Refer to caption Refer to caption Refer to caption
@@ 4.1 Embedding matrix
Construction of the embedding matrix. For a vocabulary with $n$ non-interchangeable tokens and $m$ interchangeable tokens, ${\bm{L}}\in{\mathbb{R}}^{n\times d_{\alpha}}$ represents the matrix of learnable embeddings for non-interchangeable tokens, ${\bm{\alpha}}\in{\mathbb{R}}^{1\times d_{\alpha}}$ the shared learnable embedding for interchangeable tokens, and ${\bm{\beta}}_{i}\in{\mathbb{R}}^{1\times d_{\beta}}$ the randomly-generated embedding for the $i$th interchangeable token where $1\leq i\leq m$. Note that ${\bm{\alpha}}$ and ${\bm{\beta}}_{i}$ are row vectors. A zero matrix of size $i\times j$ is represented by $\bm{0}^{i,j}$. In addition, we define two row-based L2 normalization functions $f_{bn}({\bm{X}})$ and $f_{fn}({\bm{X}})$ that divide each row ${\bm{X}}_{i,:}$ by its L2 norm $\|{\bm{X}}_{i,:}\|$. These two functions are identical but can be disabled independently from each other, hence the separation. Finally, the overall structure of the embedding matrix ${\bm{U}}$ is shown in Equation 1. In this construction, the interchangeable tokens are assumed to come after the non-interchangeable tokens. Note that it’s also possible to implement multiple sets of different interchangeable tokens via a trivial extension.
@@
Normalization. There are several concerns that warrant the heavy use of normalization while constructing ${\bm{U}}$, as seen in Equation 1. Firstly, $d_{\alpha}$ dimensions and $d_{\beta}$ dimensions should not overwhelm each other in terms of magnitude. Normalizing ${\bm{\alpha}}$ and ${\bm{\beta}}_{i}$ separately addresses this issue. The magnitude of the concatenated embedding is another concern, which is handled by the final normalization. The normalization of ${\bm{L}}$ is redundant (since the final normalization does the same operation after the concatenation with zeros) but kept in Equation 1 for readability.
@@ 4.2 Random embedding generation
Uniqueness constraint. In the normal distribution method, we don’t have any additional constraints to ensure distinguishability between vectors. However, in other two methods, we need to make sure that each interchangeable token gets assigned to a unique vector since the sampling set is finite. To achieve this quickly and space-efficiently, we define a mapping from integers to possible vectors. The unique vectors are generated by sampling $m$ unique random integers (which can be calculated efficiently using the reservoir sampling technique), and then using the defined mapping to convert these integers to the vectors. This strategy avoids materializing the whole set of possible vectors. In the hypercube vertices method, we map the binary digits of an integer in $[0,2^{d_{\beta}})$ to $\{-1,1\}$. Although “Neighboring Points” is simply the ternary version of the same idea, avoiding the zero vector requires special care. The zero vector maps to the integer $i_{z}=(3^{d_{\beta}}-1)/2$. Therefore, we define our domain as the integers in $[0,3^{d_{\beta}}-1)$ and add $1$ to the integer $i$ before converting it if $i\geq i_{z}$. Integer mapping approach for generating unique vectors works well for up to 32 dimensions, after which the limits of integer representation become an issue for reservoir sampling. Therefore, in such cases, we simply disable the uniqueness check because the exponentially growing size of the sampling set renders the probability of drawing the same sample negligible.
@@ 4.3 Projection
Weight tying. In a traditional language modeling setting, since both the embedding and projection matrices are entirely composed of learnable parameters, it’s not necessary to share them, even though there are many advantages of weight tying (Press & Wolf, 2016). However, we construct the embedding matrix manually in our method, which makes weight tying a requirement. Furthermore, since we perform our experiments on an encoder-decoder architecture in this paper, we utilize a three-way weight tying approach, whereby the embedding matrices of encoder and decoder are tied in addition to the final projection matrix. Three-way weight tying is particularly appropriate for the LTL solving task since many tokens are shared between the LTL formulae and traces.
@@
Feature normalization. Given the output of the last layer before the final projection ${\bm{v}}$ (henceforth called feature vector), instead of directly applying the final projection as in ${\bm{U}}{\bm{v}}$, we apply L2 normalization to the feature vector ${\bm{v}}$ before passing it through the final projection: ${\bm{U}}f_{fn}({\bm{v}})$. This matrix multiplication constitutes taking a dot product with each row. Since ${\bm{a}}\cdot{\bm{b}}=\|{\bm{a}}\|\|{\bm{b}}\|cos(\theta)$ where $\theta$ is the angle between ${\bm{a}}$ and ${\bm{b}}$, normalizing both the embeddings and the feature vector leaves only the cosine term to determine the logits. This forces the model to distinguish between tokens based solely on the directions, which may improve the gradient flow.
@@
Cosine loss. If we normalize both the embeddings and the feature vector, the only thing that determines each logit is the cosine of the angle between the feature vector and the embedding. Applying the softmax loss to such logits is known as cosine loss in the literature. Although cosine-based loss functions were successful in face recognition (Ranjan et al., 2017; Wang et al., 2017), it proved sensitive to hyperparameter settings in these losses. To avoid this problem, we use AdaCos loss function (Zhang et al., 2019) that scales the logits adaptively throughout training.
@@
Despite the attractiveness of AdaCos in this context, it is not directly applicable in a language modeling setting due to the additional sequence length dimension, and no prior work explored this application to the best of our knowledge. To overcome this, we modify the AdaCos loss function as follows: First, we combine the batch and length dimensions while ignoring the padding tokens, effectively treating both dimensions as batch dimensions. However, since this change greatly increases the number of batch dimensions, it can lead to numerical issues, even with the log-sum-exp trick. Therefore, we clip the scale value calculated by AdaCos to a maximum of 100 to avoid numerical issues. This loss formulation can also be used with conventional embeddings, as we do in our experiments.
Figure 2: Two annotated heatmaps visualizing the test-set edit distance between prediction and ground truth in copying task with extendable vocabulary. Both heatmaps share the same y-axis. The green box represents the number of unique characters (y-axis) and the maximum length (x-axis) in the training dataset. Each point shows the average test error, except the lower triangular part of each heatmap (gray hatch pattern) corresponding to the impossible combinations of length and unique character counts. The traditional approach (left), using ubiquitously utilized fixed (learned) token embeddings, cannot extrapolate to vocabulary expansions. The proposed method (right) enables generalization to larger vocabulary sizes at longer sequence lengths, compared to what is observed during training.